OpenStudy (anonymous):

Assume the annual number of skin cancer deaths increases geometrically. Use the starting number of 9,300 in 2000 and 200,000 in 2050. Select the number closest to the ratio of number of deaths from year to year. a) 1.0633 b) 1.3259 c) 1.6320 d) 1.0655

OpenStudy (whpalmer4):

This is just saying that \[9300*a^{50} = 200000\]and asking you to find the value of \(a\) I'm suspicious that one of the answer choices has a typo...

OpenStudy (anonymous):

Oops! I did mess up on the first answer choice! I'm sorry!

OpenStudy (whpalmer4):

Actually, I was looking at a different one!

OpenStudy (whpalmer4):

Do you know how to solve for \(a\) in that equation?

OpenStudy (anonymous):

Yes! I got a! Is that correct??:)))

OpenStudy (whpalmer4):

Yes, \(1.0633^50 = 21.5174\) and \(21.5174*9300 = 200112\) so that appears to be correct!

OpenStudy (whpalmer4):

Oops, I forgot the { } around the exponent. Should have been \[1.0633^{50} = 21.5174\]

OpenStudy (anonymous):

Okay, so, 2 questions later, it says "Compare the total number of deaths due to skin cancer expected between 2000 and 2050, if the progression were geometric and if the progression were arithmetic." Is there a formula for that that i forgot to take note of? The question means like every single year added up, for 50 years. D:

OpenStudy (anonymous):

Yay, thank you!

OpenStudy (anonymous):

Like, i know the formulas for just one year, specifically, but, gahh, 50 together??

OpenStudy (whpalmer4):

It's not entirely clear to me what the arithmetic progression would be, I'm sorry to say But the sum of an arithmetic progression isn't difficult to find. And no, you don't have to add up each term by hand :-)

OpenStudy (whpalmer4):

If you have an arithmetic sequence \(a_1,a_2,a_3,...\), and you want to add the first \(n\) terms, it's just \[S_n = \frac{n(a_n+a_1)}{2}\]

OpenStudy (anonymous):

What are 'an' and 'a1' again? i need to make sure i plugged it in right, 'cause my answer is really close to one listed but not quite..

OpenStudy (whpalmer4):

\(a_1\) is the first term \(a_n\) is the \(n\)th term

OpenStudy (anonymous):

Yeah, i'm getting 5,232,500 by plugging in 50 for n, 200,000 for the 50th term, and 9300 for the first term. The closest choice is 5,223,200. Am i doing something wrong? :(

OpenStudy (anonymous):

@whpalmer4 :(

OpenStudy (whpalmer4):

Here's a hint: why don't you subtract your answer from the closest choice (or the other way around).

OpenStudy (anonymous):

i meat to post this last night but my internet turns off at midnight! After subtracting my answer from the closest choice, i got the starting amount! So, i picked that answer. :) haha Thank you soooo sosososososo much for your help!!